OpenStudy (anonymous):

A rectangle is 5 times as long as it is wide. the perimeter is 50 cm. Find the dimensions of the rectangle. Round to the nearest tenth is necessary. Can you just tell me how to work out the problem please.

OpenStudy (anonymous):

Let width of the rectangle be x cm then length of the rectangle is 5x cm. Now, 2(5x+x)=50

OpenStudy (anonymous):

Are u following?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Now, solve for x

OpenStudy (anonymous):

well i think, 2(5x+x)=50 becomes 10x+2x=50 then 12x=50 Then 50/12? which is 4.17?

OpenStudy (anonymous):

Very correct...... U round it nearest tenth.

OpenStudy (anonymous):

4.2?

OpenStudy (anonymous):

yep... Now, width=x=4.2 cm And length =2x=2*4.2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yess, thanks sooo much! :)

OpenStudy (anonymous):

Sorry length =5x=5*4.2